How can we evaluate the change in density between the given pressure when melting point of tin at pressure of 1 bar is 505K, but increases to 508.4K at 1000 bar?
please tel the steps and formulas used.
1
Expert's answer
2010-12-20T16:22:22-0500
According the formula
PV/T = const
So, we get
103x103 Pa *V1 / 505K = 103x106 Pa *V2 / 508.4 K
But the mass of tin is const, so density is greater inV1/V2.
V1/V2 = 103,000/505 * 508.4 / 103,000,00 = 1.006*10-3
Answer: it’s greater in 1.006*10-3